Please show your work in order for me to understand problems, thank you...
1. x^2 + 11x + 24=0 Solve and show work
2. 3x^2 + 8x + 5 Factor and show work
3. (15 (21/(5 (7 Simplify radical notation
Thanks again.

Answer by edjones(8007)

1. x^2+11x+24=0
(x+8)(x+3)=0
x=-8, x=-3
.
Ed